OpenStudy (anonymous):

-5x+12y=20, x-2y=-6 solve by using the elimination method

OpenStudy (anonymous):

5x+12y=20, x-2y=-6 solve by using the elimination method

OpenStudy (anonymous):

-5x not 5x sorry!

OpenStudy (anonymous):

try it... do what i showed u last time.

OpenStudy (anonymous):

Alrighty I can solve the the rest of those on my own but now will you help me with 4 more?

OpenStudy (anonymous):

sure... post it and lemme know :)

OpenStudy (anonymous):

5x+7y=-2, 2x-7y=9

OpenStudy (anonymous):

it says you may use either the substitution method or the elimination method.

OpenStudy (anonymous):

is this a new question?

OpenStudy (anonymous):

Yes! solve by using either method